OpenStudy (anonymous):

Part 1: Determine the geometric partial sum the sum of 3 times 2 to the I minus 1 power from 1 to 7. Part 2: In complete sentences, explain the necessary steps required to find the partial sum. Include an explanation of whether this partial sum is divergent or convergent in nature.

OpenStudy (anonymous):

\[\sum_{i=1}^{7}3(2)^{i-1}\]

OpenStudy (ranga):

The 3 can be taken out of the sigma. Sum = 3(2^0 + 2^1 + 2^2 +.... + 2^6) use the formula for the geometric sum. Or in this case, since there are only 7 terms you can write each one out and add them up.

OpenStudy (ranga):

Sum = 3(1 + 2 + 4 + 8 + 16 + 32 + 64) = 3(127) = 381

OpenStudy (anonymous):

I didn't mean to close this!!!! Please don't leave @ranga

OpenStudy (ranga):

Or, using the geometric sum formula geometric sum = a(1 - r^n) / (1 - r) a = 1 r = 2 n = 7 1*(1 - 2^7) / (1 - 2) = 127. Multiply by 3 that was taken out of the sigma: 3 * 127 = 381.

OpenStudy (anonymous):

so the sum is 381?

OpenStudy (ranga):

yes.

OpenStudy (anonymous):

Thank you sooooooooooo much!!!!!!!!

OpenStudy (ranga):

you are welcome.